A motorcyclist is making an electric vest that, when connected to the motorcycle's 12 V battery, will warm her on cold rides. She is using 0.25-mm-diameter copper wire, and she wants a current of 5.5A in the wire. What length wire must she use?
Answer #1

The resistance can be calculated by the formula,

$$ R=\frac{\rho l}{A} $$

Here, \(l\) is the length of the wire, \(\rho\) is the resistivity and \(A\) is the area.

The area of the copper wire will be,

$$ \begin{aligned} A &=\pi r^{2} \\ &=\pi\left(\left(\frac{0.25 \mathrm{~mm}}{2}\right)\left(\frac{10^{-3} \mathrm{~m}}{1 \mathrm{~mm}}\right)\right)^{2} \\ &=4.91 \times 10^{-8} \mathrm{~m}^{2} \end{aligned} $$

The resistivity of copper wire is, \(\rho=1.68 \times 10^{-8} \Omega \cdot \mathrm{m}\) Also, with the help of Ohm's law, \(R=\frac{V}{I}\)

Substitute \(\frac{V}{I}\) for \(R\) in equation (1).

$$ \begin{aligned} \frac{V}{I} &=\frac{\rho l}{A} \\ l &=\frac{V A}{I \rho} \\ &=\frac{(12 \mathrm{~V})\left(4.91 \times 10^{-8} \mathrm{~m}^{2}\right)}{(5.5 \mathrm{~A})\left(1.68 \times 10^{-8} \Omega \cdot \mathrm{m}\right)} \\ &=6.4 \mathrm{~m} \end{aligned} $$